# Model order reduction via Lie groups

### v1

@@ Abstract
Lie groups and their actions are ubiquitous in the description of physical systems, and we explore implications in the setting of model order reduction (MOR). We present a novel framework of MOR via Lie groups, called MORLie, in which high-dimensional dynamical systems on manifolds are approximated by low-dimensional dynamical systems on Lie groups. In comparison to other Lie group methods we are able to attack non-equivariant dynamics, which are frequent in practical applications, and we provide new non-intrusive MOR methods based on the presented geometric formulation. We also highlight numerically that MORLie has a lower error bound than the Kolmogorov $N$-width, which limits linear-subspace methods. The method is applied to various examples: 1. MOR of a simplified deforming body modeled by a noisy point cloud data following a sheering motion, where MORLie outperforms a naive POD approach in terms of accuracy and dimensionality reduction. 2. Reconstructing liver motion during respiration with data from edge detection in ultrasound scans, where MORLie reaches performance approaching the state of the art, while reducing the training time from hours on a computing cluster to minutes on a mobile workstation. 3. An analytic example showing that the method of freezing (a previous Lie group method for MOR introduced in [1]) is analytically recovered as a special case, showing the generality of the geometric framework.
@@ 1 Introduction
Figure 1: Summary of MORLie: Given a full order model (FOM) based on a vector field $X\in\Gamma(T\mathcal{M})$, a reduced order model (ROM) on a lower-dimensional Lie group $G$ is expressed using an action $\Phi:G\times\mathcal{M}\rightarrow\mathcal{M}$ and a reduction map $\rho:\mathcal{M}\rightarrow\mathfrak{g}$. The goal is to choose $(G,\Phi,\rho)$ such that $\bar{x}(t)=\Phi(g(t),x_{0})$ approximately follows the FOM dynamics $\dot{\bar{x}}\approx X(\bar{x})$.
@@
We show that our method induces a distribution on the state-space manifold, rather than picking a submanifold of it. Thus initial states $u_{0}\in\mathcal{M}$ are not projected to a submanifold, and approximations can reproduce initial states exactly.
@@
    Multiple example applications: point clouds undergoing sheering motions, liver tracking during respiration, and analytically recovering the method of freezing [1] (Section 6).
The article is ordered as follows. We conclude the introduction with related literature and notation. Section 2 provides background on Lie groups and their actions on manifolds. Section 3 introduces the main problem, i.e., to describe an MOR method that results in a ROM on a Lie group and to conceptually connect such a method to existing theory on MOR on manifolds. Next, Section 4 formally describes MORLie and introduces the group Kolmogorov $N$-width. Afterwards, Section 5 presents intrusive and non-intrusive optimization methods, and initial ideas for hyperreduction. Section 6 provides example applications and a discussion of the results. A conclusion and outlook are given in Section 7.
@@ 1.1 Related literature
Submanifold methods for MOR are formalized in a differential geometric language in [25]. Most nonlinear approaches to MOR can be seen as particular cases of submanifold methods: transformation-based methods [33, 34, 35, 36] apply linear or nonlinear, time-dependent transformations to a fixed and reduced basis (still allowing for time-varying coefficients). A similar formalism can also be derived by transforming snapshot matrices with nonlinear transformations [37], which can reduce the Kolmogorov $N$-width. By transforming a reduced basis, such methods effectively represent the system on a submanifold of the state-space. In [38] integration on Lie groups $SO(n),SP(n)$ is used for geometrically exact integration of time-varying reduced bases for Hamiltonian systems. We interpret this as a submanifold method, since the Lie groups are made to act on a Stiefel (sub-)manifold [39] that the time-varying reduced basis of the total system.
@@ 2.1 Lie groups and their actions
In other words, the infinitesimal generator $X_{\widetilde{A}}(x)$ describes the velocity of a curve $g(t)\cdot x$ on $\mathcal{M}$ at $t=0$, if $g(0)=e$ and $\dot{g}(0)=\widetilde{A}$.
@@
The first statement is a direct consequence of [53, Proposition 7.26]. Integrability of $\Delta$ follows from [53, Proposition 19.2] and Frobenius’ Theorem [53, Theorem 19.12]. Finally, tangency of $\Delta_{x}$ and $\mathcal{O}(x)$ is a direct consequence of Defs. 2.3, 2.5 and 2.7.
@@ 2.2 The Kolmogorov N-width
In the context of MOR the Kolmogorov $N$-width is analyzed for the set of solution snapshots collected in $S$, and it provides a lower bound on the worst-case error of linear-subspace methods in approximating $S$. Analytic estimates of the Kolmogorov $N$-width are well-known for certain problem classes: for example, the Kolmogorov $N$-width is known to decay exponentially for diffusion problems [54], but only slowly (order $\sim\frac{1}{\sqrt{N}}$) for certain linear transport and wave problems [55]. While achieving the theoretical best case given by $d_{N}(S)$ is difficult, it can often be approached. Yet, the slow decay for linear transport suggests that linear-subspace methods are ill-suited for most transport dominated problems [15]. For recent nonlinear alternatives to the Kolmogorov $N$-width, for which the linear-subspace Kolmogorov $N$-width generally presents a lower bound, refer also to [56, 57].
@@ 3 Problem formulation
    How does this approach relate to submanifold methods? In Appendix C.
@@ 4.1 ROM dynamics
We want to choose the Lie group dynamics such that the evolution of $\Phi(g_{\mu}(t),x_{\mu,0})$ closely follows (w.r.t. a to-be-defined metric) the full-order state $x_{\mu}(t)$. Since the FOM dynamics $X_{\mu}(x_{\mu})$ vary with $x_{\mu}$, the optimal Lie group dynamics should also vary with $x_{\mu}$. We encode this by a map $\rho_{\mu}:\mathcal{M}\rightarrow\mathfrak{g}$, and immediately state the main technical result of this subsection:
@@ 4.2 The group Kolmogorov N-width
Thus, $d_{\Xi_{N}}(S,x_{0})$ is the worst case error over the orbit $\mathcal{O}(x_{0})$, for the best choice of admissible pair $(G,\Phi)\in\Xi_{N}$. In turn, $d^{T}_{\Xi_{N}}(S)$ formalizes the error between $x_{\mu}(t)$ and $\bar{x}_{\mu}(t)$ (cf. (17)), returning the worst case over all initial conditions. Finally, $d^{T,\tau}_{\Xi_{N}}(S)$ measures a worst-case error between FOM-trajectories and ROM-trajectories over a time-horizon $\tau$, and without restricting itself to ROM-trajectories starting at $x_{\mu,0}$. Alternatively, we can interpret $\frac{1}{\tau}d^{T,\tau}_{\Xi_{N}}(S)$ as a bound on the increase of $\textnormal{dist}\big(\bar{x}(t),x(t)\big)$ per $\tau$. We note that Definitions (28) and (29) coincide for $\tau=T$ ($d^{T}_{\Xi_{N}}(S)=d^{T,T}_{\Xi_{N}}(S)$), in which case the supremum over $t\in[0,T-\tau]$ in (29) is over a single point $t=0$.
@@
Yet, affine MOR only provides a marginal advantage over linear MOR since $\Xi^{\textnormal{Lin}}_{N+1}$ includes $\textnormal{Aff}_{N}$, such that $d_{\Xi^{\textnormal{Lin}}_{N+1}}(S,0)\leq d_{\Xi^{\textnormal{Aff}}_{N}}(S,0)$. Section 6 will show more significant examples of breaking the Kolmogorov barrier by means of group methods, and also provide estimates for specific group Kolmogorov $N$-widths.
@@ 5 Optimization in MORLie
We want to choose $(G,\Phi,\rho_{\mu})$ to arrive at ROM dynamics that can be evaluated more efficiently than the FOM dynamics, and enable reconstructed solutions to approximate FOM solutions (cf. (17)). To this end, intrusive and non-intrusive methods for MORLie are presented, also preparing for the examples in Section 6. In both cases we cast the process of determining $(G,\Phi,\rho_{\mu})$ for a given $(\mathcal{M},X_{\mu},S)$ into the form of an optimization problem.
@@ 5.3 Intrusive and non-intrusive MORLie
For a first intrusive cost-function, we consider a smooth Riemannian manifold $\mathcal{M}$ with metric-induced norm $\|\cdot\|$ on local tangent-spaces $T_{x}\mathcal{M}$, and metric-induced volume form. Then we define
@@
$=\sum_{\mu_{i}\in\mathcal{P}}\int_{\mathcal{M}}\|X_{\mu_{i}}(x)-X_{\rho_{\theta,\mu_{i}}(x)}(x)\|\,.$
@@
where $\Pi_{\Delta}:T\mathcal{M}\rightarrow\Delta$ is the metric-projection onto the induced distribution $\Delta$ (cf. Def. 2.7) and ${X_{\mathfrak{g}}}^{\dagger}:\Delta\rightarrow\mathfrak{g}$ inverts the infinitesimal generator ($\forall{\widetilde{A}\in\mathfrak{g}}:{X_{\mathfrak{g}}}^{\dagger}X_{\widetilde{A}}=\widetilde{A}$) [^1] The operator ${X_{\mathfrak{g}}}^{\dagger}\Pi_{\Delta}$ is a coordinate-free version of a Moore-Penrose pseudo-inverse of $X_{\mathfrak{g}}$.. Further, if $X_{\mu}$ is smooth, then so is $\rho^{*}_{\mu}$.
@@
For a discrete implementation consider state snapshots $S$ collecting discrete trajectories $x_{i,k}:=x_{\mu_{i}}(t_{i,k})$ and parameters $\mu_{i}$:
@@
For a velocity-free non-intrusive cost-function we consider a smooth space-time manifold $\mathcal{M}\times\mathbb{R}$ with spatial metric $\textnormal{dist}:\mathcal{M}\times\mathcal{M}\rightarrow\mathbb{R}_{+}$, and state-time snapshots $S$ collecting discrete trajectories $x_{i,k}:=x_{\mu_{i}}(t_{i,k})$, time-instances $t_{i,k}$ and parameters $\mu_{i}$:
@@
We note that the cost (52) is strongly related to (29) for small $\tau=\Delta t_{i,k}$, whereas the velocity based cost (50) may be seen as a limiting case as $\lim_{\tau\rightarrow 0}$.

### v2

@@ Abstract
Lie groups and their actions are ubiquitous in the description of physical systems, and we explore implications in the setting of model order reduction (MOR). We present a novel framework of MOR via Lie groups, called MORLie, in which high-dimensional dynamical systems on manifolds are approximated by low-dimensional dynamical systems on Lie groups. In comparison to other Lie group methods we are able to attack non-equivariant dynamics, which are frequent in practical applications, and we provide new non-intrusive MOR methods based on the presented geometric formulation. We also highlight numerically that MORLie has a lower error bound than the Kolmogorov $N$-width, which limits linear-subspace methods. The method is applied to various examples: 1. MOR of a simplified deforming body modeled by noisy point cloud data following a sheering motion, where MORLie outperforms a naive POD approach in terms of accuracy and dimensionality reduction. 2. Reconstructing liver motion during respiration with data from edge detection in MRI scans, where MORLie reaches performance approaching the state of the art, while reducing the training time from hours on a computing cluster to minutes on a mobile workstation. 3. An analytic example showing that the method of freezing (a previous Lie group method for MOR introduced in [1]) is analytically recovered as a special case, showing the generality of the geometric framework.
@@ 1 Introduction
Figure 1: Summary of MORLie. Given a full order model (FOM) defined by a vector field $X$ on a state-manifold $\mathcal{M}$, we construct a reduced order model (ROM) on a lower-dimensional Lie group $G$. A group action $\Phi:G\times\mathcal{M}\rightarrow\mathcal{M}$ is used to reconstruct states via $\bar{x}(t)=\Phi(g(t),x_{0})$, where the evolution of $g(t)$ is determined by a reduced vector field $\rho:\mathcal{M}\rightarrow\mathfrak{g}$. The ROM dynamics on $G$ are defined as $\dot{g}={R_{g}}_{*}\rho(\bar{x})$, and induce approximated dynamics on $\mathcal{M}$ that are constrained to group orbits. The goal is to choose $(G,\Phi,\rho)$, such that the reconstructed solutions $\bar{x}(t)$ closely follows the FOM dynamics $\dot{\bar{x}}\approx X(\bar{x})$.
@@
While submanifold methods for MOR restrict trajectories to lie on a fixed submanifold, we show that (3) instead restricts the dynamics to directions tangent to group orbits. Thus, rather than constraining where states can lie, we constrain how they are allowed to move. In this way, initial conditions are preserved: for any initial state $u_{0}\in\mathcal{M}$, the approximation evolves on the orbit through $u_{0}$, so no projection onto a reduced manifold is required. More formally, our method induces a low-dimensional distribution on the state-space manifold, rather than selecting a submanifold of it.
@@
    Multiple example applications: analytically recovering the method of freezing [1] (Section 6), point clouds undergoing sheering motions and liver tracking during respiration (Section 7).
The article is ordered as follows. We conclude the introduction with related literature and notation. Section 2 provides background on Lie groups and their actions on manifolds. Section 3 introduces the main problem, i.e., to describe an MOR method that results in a ROM on a Lie group and to conceptually connect such a method to existing theory on MOR on manifolds. Next, Section 4 formally describes MORLie and introduces the group Kolmogorov $N$-width. Afterwards, Section 5 presents intrusive and non-intrusive optimization methods, and initial ideas for hyperreduction. Section 6 provides applications of MORLie to linear transport and the method of freezing [1]. Section 7 provides numerical example applications and a discussion of the results. A conclusion and outlook are given in Section 8.
@@ 1.1 Related literature
Submanifold methods for MOR are formalized in a differential geometric language in [25]. Most nonlinear approaches to MOR can be seen as particular cases of submanifold methods: transformation-based methods [33, 34, 35, 36] apply linear or nonlinear, time-dependent transformations to a fixed and reduced basis (still allowing for time-varying coefficients). A similar formalism can also be derived by transforming snapshot matrices with nonlinear transformations [37], which can reduce the Kolmogorov $N$-width. By transforming a reduced basis, such methods effectively represent the system on a submanifold of the state-space. The work of [38, 39] uses a shifted POD, where the shift can be seen as a particular choice of Lie group action – we generalize towards accommodating arbitrary Lie group actions, see e.g., Section 6.2. In [40] integration on Lie groups $SO(n),SP(n)$ is used for geometrically exact integration of time-varying reduced bases for Hamiltonian systems. We interpret this as a submanifold method, since the Lie groups are made to act on a Stiefel (sub-)manifold [41] that the time-varying reduced basis of the total system.
@@
For a review of the Kolmogorov $N$-width and recent nonlinear widths extending the concept, see the dedicated Section 2.2.
@@ 2.1 Lie groups and their actions
Intuitively, the infinitesimal generator $X_{\widetilde{A}}(x)$ describes the velocity at $x$ generated by the action of $\exp(\widetilde{A}t)$.
@@
In the present work, the relevant distribution will be the span of infinitesimal generators.
@@
The first statement is a direct consequence of [55, Proposition 7.26]. Integrability of $\Delta$ follows from [55, Proposition 19.2] and Frobenius’ Theorem [55, Theorem 19.12]. Finally, tangency of $\Delta_{x}$ and $\mathcal{O}(x)$ is a direct consequence of Defs. 2.3, 2.5 and 2.7.
@@ 2.2 The Kolmogorov N-width
In the context of MOR the Kolmogorov $N$-width is analyzed for the set of solution snapshots collected in $S$, and it provides a lower bound on the worst-case error of linear-subspace methods in approximating $S$. Analytic estimates of the Kolmogorov $N$-width are well-known for certain problem classes: for example, the Kolmogorov $N$-width is known to decay exponentially for diffusion problems [56], but only slowly (order $\sim\frac{1}{\sqrt{N}}$) for certain linear transport and wave problems [57]. While achieving the theoretical best case given by $d_{N}(S)$ is difficult, it can often be approached. Yet, the slow decay for linear transport suggests that linear-subspace methods are ill-suited for most transport dominated problems [15].
Lower bounds for nonlinear MOR methods are given by nonlinear alternatives to the linear-subspace Kolmogorov $N$-width. For these nonlinear widths the linear width (14) generally presents an upper bound, refer also to [58, 59].
Denoting $D_{N}:\mathbb{R}^{N}\rightarrow\mathcal{M}$ a continuous, Lipschitz decoder function, constant $\gamma>0$, then the Lipschitz $N$-width [60] of $S\subset\mathcal{M}$ is
$d^{\textnormal{Lip}}_{N}(S)=$ $\inf_{\begin{subarray}{c}D_{N},\|\cdot\|\\ \gamma\textnormal{-Lipschitz}\end{subarray}}\sup_{x\in S}\inf_{\|\bar{g}\|\leq B}\|x-D_{N}(\bar{g})\|\,$, (15)
where the final infimum is also over a suitably restricted class of norms $\|\cdot\|$ on $\mathcal{M}$.
In [61] the Sectional Komogorov $N$-width is presented. Denote by $\sigma:\mathcal{P}\rightarrow\mathcal{M}$, $\sigma\in\Gamma$ a section of a fiber bundle with the parameter manifold $\mathcal{P}$ as the base space, and the state manifold $\mathcal{M}$ as the fiber. Then the linear Sectional $N$-width measures distance to an optimally chosen linear sub-bundle $\Gamma_{N}$:
$d^{\textnormal{Sec}}_{N}(S)=\inf_{\begin{subarray}{c}\Gamma_{N}\subset\Gamma\;\\ \dim\Gamma_{N}\leq N\\ \textnormal{linear}\end{subarray}}\sup_{\mu\in\mathcal{P}}\inf_{\sigma_{N}\in\Gamma_{N}}\|\sigma(\mu)-\sigma_{N}(\mu)\|\,.$ (16)
Non-linear extensions in the same work [61] emphasize that $\mathcal{M}$ can be a manifold, in which case $\Gamma_{N}$ is a sub-bundle for which the fiber $\sigma_{N}(\mu)$ is a submanifold of $\sigma(\mu)$.
@@ 3 Problem formulation
We will mention it explicitly when finite or countably infinite snapshot sets $S$ are considered, or when $S\subset\mathcal{M}\times\mathcal{P}\times[0,T]$ by abuse of notation.
@@
    How does this approach relate to submanifold methods? In Appendix D.
@@ 4.1 ROM dynamics
We want to choose the Lie group dynamics such that the evolution of $\Phi(g_{\mu}(t),x_{\mu,0})$ closely follows (w.r.t. a to-be-defined metric) the full-order state $x_{\mu}(t)$. Since the FOM dynamics $X_{\mu}(x_{\mu})$ vary with $x_{\mu}$, the optimal Lie group dynamics should also vary with $x_{\mu}$. We encode this by a map $\rho_{\mu}:\mathcal{M}\rightarrow\mathfrak{g}$, and immediately state the main technical result of this subsection: the induced dynamics on $\mathcal{M}$ generated by dynamics on $G$. The key idea is that a vector in the Lie algebra produces a vector field on $\mathcal{M}$ via the infinitesimal generator.
@@
Thus, $\rho$ defines dynamics on $G$, which induce approximated dynamics $\bar{X}$ on $\mathcal{M}$, constrained to the orbits $\mathcal{O}(x)$ and tangent to the distribution $\Delta$.
@@ 4.2 The group Kolmogorov N-width
Intuitively, the group width measures how well solution trajectories can be approximated by a single group orbit through the initial state.
In particular, $d_{\Xi_{N}}(S,x_{0})$ is the worst case error over the orbit $\mathcal{O}(x_{0})$, for the best choice of admissible pair $(G,\Phi)\in\Xi_{N}$. In turn, $d^{T}_{\Xi_{N}}(S)$ formalizes the error between FOM-trajectories $x_{\mu}(t)$ and ROM-trajectories $\bar{x}_{\mu}(t)$ (cf. (19)), returning the worst case over all initial conditions. Finally, $d^{T,\tau}_{\Xi_{N}}(S)$ measures a worst-case error between $x_{\mu}(t)$ and $\bar{x}_{\mu}(t)$ over a time-horizon $\tau$, and without restricting itself to ROM-trajectories starting at $x_{\mu,0}$. Alternatively, we can interpret $\frac{1}{\tau}d^{T,\tau}_{\Xi_{N}}(S)$ as a bound on the increase of $\textnormal{dist}\big(\bar{x}(t),x(t)\big)$ per $\tau$. We note that Definitions (30) and (31) coincide for $\tau=T$ ($d^{T}_{\Xi_{N}}(S)=d^{T,T}_{\Xi_{N}}(S)$), in which case the supremum over $t\in[0,T-\tau]$ in (31) is over a single point $t=0$.
@@
Yet, affine MOR only provides a marginal advantage over linear MOR since $\Xi^{\textnormal{Lin}}_{N+1}$ includes $\textnormal{Aff}_{N}$, such that $d_{\Xi^{\textnormal{Lin}}_{N+1}}(S,0)\leq d_{\Xi^{\textnormal{Aff}}_{N}}(S,0)$. Sections 6, 7 will show more significant examples of breaking the Kolmogorov barrier by means of group methods, and Section 6 provides a specific group Kolmogorov $N$-width of linear transport.
@@
The Lipschitz $N$-width (15) can also be recovered from (29):
Proposition 4.7
Assume that $\mathcal{M}$ is a Hilbert space with $\textnormal{dist}(x,y)=\|x-y\|$, $T^{N}=S^{1}\times\cdots\times S^{1}$ is the $N$-torus. Let $D_{N}:\mathbb{R}^{N}\rightarrow\mathcal{M}$ denote the decoder function in (15), with constant $\gamma>0$ and $\|\cdot\|$ from a suitably restricted class of norms on $\mathcal{M}$. Also construct a discontinuous embedding $\textnormal{Emb}:T^{N}\rightarrow\{x\in\mathbb{R}^{N}\;|\;\|x\|\leq B\}$ mapping $N$ copies of $S^{1}$ to $N$ copies of the half-open interval $(-B,B]\in\mathbb{R}$. Define
$\Xi^{\textnormal{Lip}}_{N}:=\{$ $(T^{N},\Phi)\;|\;\Phi(g,x):=x+D_{N}(\textnormal{Emb}(g))\,$, (43)
$D_{N}\textnormal{is }\gamma\textnormal{-Lipschitz}\}\,.$
Then we recover the Lipschitz $N$-width (15) from (29)
$d^{\textnormal{Lip}}_{N}(S)=\inf_{\|\cdot\|_{\mathcal{M}}}d_{\Xi^{\textnormal{Lip}}_{N}}(S,0)\,.$ (44)
Proof 4.7.1
The orbit $\mathcal{O}(0)$ is
$\mathcal{O}_{(T^{N},\Phi)}(0)=D_{N}(\textnormal{Emb}(T^{N}))\subset\mathcal{M}\,.$ (45)
Hence, the first infimum in $d_{\Xi^{\textnormal{Lip}}_{N}}(S,0)$ is
$\inf_{y\in\mathcal{O}(0)}\|x-y\|$ $=\inf_{g\in T^{N}}\|x-D_{N}(\textnormal{Emb}(g))\|$ (46)
$=\inf_{\|\bar{g}\|\leq B}\|x-D_{N}(\bar{g})\|\,$, (47)
where the final step holds since the image of Emb is dense in the closed ball $\{x\;|\;\|x\|\leq B\}$, and we recover
$\inf_{\|\cdot\|}d_{\Xi^{\textnormal{Lip}}_{N}}(S,0)$ $=\inf_{\begin{subarray}{c}(G,\Phi)\in\Xi^{\textnormal{Lip}}_{N}\\ \|\cdot\|\end{subarray}}\;\sup_{x\in S}\;\inf_{y\in\mathcal{O}(0)}\textnormal{dist}(x,y)$ (48)
$=\inf_{\begin{subarray}{c}D_{N},\|\cdot\|\\ \gamma\textnormal{-Lipschitz}\end{subarray}}\;\sup_{x\in S}\inf_{\|\bar{g}\|\leq B}\|x-D_{N}(\bar{g})\|\,$, (49)
which is equivalent to (15).
Here, the Lie group formalism is taken very loosely: the map $\Phi(x,g)=x+D_{N}(\textnormal{Emb}(g))$ in (43) is not explicitly restricted to be a homomorphism (cf. Def. 2.1), with the result that properties of the orbits (cf. Def. 2.4) and induced distribution (cf. Def. 2.8) are no longer guaranteed. The resulting orbits may self-intersect, thus overparameterizing $\mathcal{M}$ – this problem would be avoided if the homomorphism property can be enforced.
A ROM in MORLie can also be evaluated by means of the Sectional $N$-width (16).
Proposition 4.8
Assume that $\mathcal{M}$ is a Hilbert space with $\textnormal{dist}(x,y)=\|x-y\|$. Let $\Phi_{\mu}(g,x)$ be a parameter-dependent action. Define
$\Xi^{\textnormal{Sec}}_{N}:=\{$ $(\mathbb{R}^{N},\Phi_{\mu})\;|\;\Phi_{\mu}(g,x):=x+g^{i}e_{\mu,i}\,$, (50)
$\{e_{\mu,1},\cdots,e_{\mu,N}\}\subseteq\mathcal{M}\textnormal{ lin. indep.}\}\,.$
Any such action $\Phi_{\mu}$ induces a linear section
$\sigma_{N}(\mu)=\mathcal{O}_{(\mathbb{R}^{N},\Phi_{\mu})}(x_{\mu})\,.$ (51)
$d^{\textnormal{Sec}}_{N}(S)=d_{\Xi^{\textnormal{Lip}}_{N}}(S,0)\,.$ (52)
I.e., a parameterized linear action gives rise to a linear section $\sigma_{N}(\mu)$ with varying linear fiber. Similarly, nonlinear sections can be induced by general parameterized actions.
@@ 5 Optimization in MORLie
We want to choose $(G,\Phi,\rho_{\mu})$ to arrive at ROM dynamics that can be evaluated more efficiently than the FOM dynamics, and enable reconstructed solutions to approximate FOM solutions (cf. (19)). To this end, intrusive and non-intrusive methods for MORLie are presented, also preparing for the examples in Sections 6, 7. In both cases we cast the process of determining $(G,\Phi,\rho_{\mu})$ for a given $(\mathcal{M},X_{\mu},S)$ into the form of an optimization problem.
@@ 5.3 Intrusive and non-intrusive MORLie
For a first intrusive cost-function, we consider a smooth Riemannian manifold $\mathcal{M}$ with metric-induced norm $\|\cdot\|$ on local tangent-spaces $T_{x}\mathcal{M}$, and metric-induced volume form. With a finite parameter set $\mathcal{P}_{f}\subset\mathcal{P}$ we define
@@
$=\sum_{\mu_{i}\in\mathcal{P}_{f}}\int_{\mathcal{M}}\|X_{\mu_{i}}(x)-X_{\rho_{\theta,\mu_{i}}(x)}(x)\|\,.$
@@
where $\Pi_{\Delta}:T\mathcal{M}\rightarrow\Delta$ is the metric-projection onto the induced distribution $\Delta$ (cf. Def. 2.7) and ${X_{\mathfrak{g}}}^{\dagger}:\Delta\rightarrow\mathfrak{g}$ inverts the infinitesimal generator ($\forall{\widetilde{A}\in\mathfrak{g}}:{X_{\mathfrak{g}}}^{\dagger}X_{\widetilde{A}}=\widetilde{A}$) [^1] The operator ${X_{\mathfrak{g}}}^{\dagger}\Pi_{\Delta}$ is a coordinate-free version of a Moore-Penrose pseudo-inverse of $X_{\mathfrak{g}}$, i.e., corresponding to algebraic inversion.. Further, if $X_{\mu}$ is smooth, then so is $\rho^{*}_{\mu}$.
@@
For a discrete implementation consider a finite set of state-parameter snapshots $S$ collecting discrete trajectories $x_{i,k}:=x_{\mu_{i}}(t_{i,k})$ and parameters $\mu_{i}$:
@@
For a velocity-free non-intrusive cost-function we consider a smooth space-time manifold $\mathcal{M}\times\mathbb{R}$ with spatial metric $\textnormal{dist}:\mathcal{M}\times\mathcal{M}\rightarrow\mathbb{R}_{+}$, and finite set of state-time-parameter snapshots $S$ collecting discrete trajectories $x_{i,k}:=x_{\mu_{i}}(t_{i,k})$, time-instances $t_{i,k}$ and parameters $\mu_{i}$:
@@
We note that the cost (64) corresponds to a discrete-time flow-matching, which is strongly related to the worst case orbit approximation (31) for small $\tau=\Delta t_{i,k}$. The velocity based cost (62) is an instantaneous projection error, and may be seen as a limiting case of (31) as $\lim_{\tau\rightarrow 0}$.